# Asymptotic and compound e-values: multiple testing and empirical Bayes

### v3

@@ Appendix F Proofs for Section 6
We can argue analogously for any $m_{N}\in\mathcal{M}_{N}(\varepsilon)$ by recalling that $H^{*}\otimes G_{K}\stackrel{{\scriptstyle d}}{{\rightsquigarrow}}H^{*}\otimes G^{*}$. Thus, by taking $K\to\infty$ and then $\varepsilon\to 0$, we have shown (32). In Lemma F.1 below, we will show that (31) implies that
@@
Lemma F.1 (NPMLE convergence).
@@
for $Z\sim\mathrm{N}(0,1)$. By Proposition 3.12 it follows that $E^{(n)}$ are asymptotic e-variables. To upgrade this result to strongly asymptotic, we can apply Theorem 2.5 of Giné et al. [1997] to show that the sequence $(E^{(n)})_{n\in\mathbb{N}}$ is uniformly integrable and then conclude the proof by Proposition 3.7.
@@
which also means that $E^{(n)}$ goes to $\infty$ with probability 1. The argument for $\tilde{E}^{(n)}$ is analogous and omitted. The argument for $\hat{\sigma}^{(n)}$ follows from Proposition 3.12. ∎
F.3 Proof of Proposition 6.6
@@
Appendix G Proofs for Section 7
G.1 Proof of Proposition 7.3
Proof.
Hence, further arguing as in the proof of Theorem 2.3, for any $t>0$,
G.2 Proof of Proposition 7.4
Proof.

### v4

@@ Appendix I Proofs for Section 6
We can argue analogously for any $m_{N}\in\mathcal{M}_{N}(\varepsilon)$ by recalling that $H^{*}\otimes G_{K}\stackrel{{\scriptstyle d}}{{\rightsquigarrow}}H^{*}\otimes G^{*}$. Thus, by taking $K\to\infty$ and then $\varepsilon\to 0$, we have shown (35). In Lemma I.1 below, we will show that (34) implies that
@@
Lemma I.1 (NPMLE convergence).
@@
for $Z\sim\mathrm{N}(0,1)$. By Example 3.11 it follows that $E^{(n)}$ are asymptotic e-variables. To upgrade this result to strongly asymptotic, we can apply Theorem 2.5 of Giné et al. [1997] to show that the sequence $(E^{(n)})_{n\in\mathbb{N}}$ is uniformly integrable and then conclude the proof by Proposition 3.7.
@@
which also means that $E^{(n)}$ goes to $\infty$ with probability 1. The argument for $\tilde{E}^{(n)}$ is analogous and omitted. The argument for $\hat{\sigma}^{(n)}$ follows from Example 3.11. ∎
I.3 Proof of Proposition 6.6